Consider point ions of mass M and charge e immersed in a uniform sea of conduction electrons. The ions are imagined to be in stable equilibrium when at regular lattice points. If one ion is displaced a small distance r from its equilibrium position, the restoring force is largely due to the electric charge within the sphere of radius r centered at the equilibrium position. Take the number of density of ions (or of conduction electrons) as 3/(4pi R^3) , which defines R.

(a) Show that the frequency of a single ion set into oscillation is  omega = sqrt{e^2/(MR^3)} .

(b) Estimate the value of this frequency for sodium, given that the atomic mass of sodium is 23g/mol, the metallic sodium concentration is n2.65 10)cm3 , and e 4.8 10-10)esu .

(c) Estimate the order of magnitude of the sound velocity in metal.

(All answers please be as brief as possible but clearly written)
Answer #1

a) The number density of ions is 3/4πR3, the charge density will be 3e/4πR3
When an ion is displaced by a small distance r from its equilibrium position,the total charge contributing to the restoring force is:
q=3e/4πR3 * 4πr3/3=er3/R3
Md2r/dt2=-qe/r2=-e2r/R3
which gives an oscillation frequency of w=sqrt(e2/MR3)
b) R=cuberoot(3/8π)*a=cuberoot(3/8π)*429.06pm = 2.1126*10-10m
M= 22.99*1:660*10-27kg = 3:818*10-26kg
e= 11*1.602*10-19C = 1.762*10-18C
w=sqrt(e2/4π*ε0*MR3)= 2.8*1014rad/s
c) The wave length to be of order of lattice constant, 1010 m, sound velocity will be
v=w/k= w*λ/2π = 104 m/s